rom finish tmes from the full Pitsburgh Marathom in 2014. Beow is o 10 finish times (hours)from the 18-29 male and 18-29 fSemaliledevisom a. Compute the (8 points) b. On the same mamber line, sketch box plots of both data sets. (& points) c. Explain similarities/differences in the data. (8 points) d. Which dataset is more e. A runner finishes at 3.93 hours. What is the z-score in both sets? In which group did the runner Simis better compared to the rest of the pack of participants based on the sample? (& points) nean, median, mode, 01, Q3, 1OR, variation and standard deviation for both sets of dats variant? Prove mathematically. (8points)
Answer #1

a) male data summary

sum= 41.41 and n=10

mean= 4.141

median=4.155

mode= 4.09 (REPEATED TWO TIMES)

Quartile Q1: 3.76

Quartile Q3: 4.54

IQR= Q3-Q1= 4.54- 3.76 = 0.78

VARIANCE= 0.392

STANDARD DEVIATION= 0.626

Female data summary

Average (mean): ?=4.369

Minimum: 3.09
Maximum: 5.21
Variance: 0.329089
Corrected sample standard deviation s=0.604693678191

Median: 4.515
Quartile Q1: 3.99
Quartile Q3: 4.76

IQR= Q3-Q1= 4.76-3.99= 0.77

MODE=ALL VALUES JUST APPEARED ONCE.

I USED EXCEL TO GET ALL VALUES.

2.99 3.78.158.54 522

FEMALE 3.09 3.99 4.54576 5.21 0

COEFFICIENT OF VARIATION OF MALE DATA SET= s1/xbar = 0.1512

COEFFIENCIENT OF VARIATION OF FEMALE DATA SET = s2/ybar= 0.138

From above we can say that variation in male data set is slightly more than female dataset.

PLEASE REPOST REST QUESTIONS ALONG WITH DATA AND DESCRIPTIVE STATISTICS.